OpenStudy (anonymous):

THE AREA OF A TRIANGLE IS 27 SQUARE FEET. ITS HEIGHT IS THREE TIMES THE LENGTH OF ITS BASE. FIND THE HEIGHT AND BASE OF THE TRIANGLE.

OpenStudy (nathan917):

hight-9 base-3 Its Base x height

jhonyy9 (jhonyy9):

no @nathan917 area of a triangle is base time height divide 3 ATTENTION !!! so you know that area = 27 ft^2 the height is 3 time the length of base so note h=3b than you can rewriting area = b*3b/3 =b*b=b^2 so this mean that 27 = b^2 than b = sqrt 27 = 3sqrt3 b=3sqrt3 h= 3b = 3*3sqrt3 = 9sqrt3 h=9sqrt3